OpenStudy (anonymous):

Convert the given function into a logarithmic function: y=14.38e^(-.36x)

OpenStudy (anonymous):

y = 14.38e^(-0.36x) ln y = ln (14.38e^(-0.36x)) = ln 14.38 + ln e^(-0.36x) = ln 14.38 - 0.36x Formulas used: ln (ab) = ln a + b ln e^n = n

OpenStudy (anonymous):

how would i graph that @TURITW